OpenStudy (anonymous):

Use binomial expansions to find the expansion in powers of x up to x^4 of (1-x^3)^6 * (1-x)^-6

ganeshie8 (ganeshie8):

\(\large (1-x^3)^6 * (1-x)^{-6} \) \(\large (1-x^3)^6 * \frac{1}{(1-x)^{6}} \) \(\large \left( \frac{1-x^3}{1-x} \right)^6 \)

ganeshie8 (ganeshie8):

\(\large \left( \frac{1^3-x^3}{1-x} \right)^6 \) \(\large \left( \frac{(1-x)(x^2+x+1)}{1-x} \right)^6 \) \(\large \left( x^2+x+1 \right)^6 \)

ganeshie8 (ganeshie8):

now try expanding

OpenStudy (anonymous):

Ok, yes I have expanded that (it's actually the same as another question I've already answered) and it comes out with the right answer. Thank you!

ganeshie8 (ganeshie8):

cool :) you did the samy way as above... or any other easy way out ?

OpenStudy (anonymous):

Same as your method

ganeshie8 (ganeshie8):

okie :) binomial expansion we cant do for negatives is it

ganeshie8 (ganeshie8):

wud the method above, we have to apply binomial theorem two times... and combine like terms which is a pain

ganeshie8 (ganeshie8):

was thinking, if there is any easy way...

ganeshie8 (ganeshie8):

nvm ! we cant do binomial expansion for negative powers